A digital clock has a resistance of 1.00 x 10^4 W and is plugged into a 110.0-V outlet. If the owner pays $0.05 per kWh, how much does it cost to operate the clock for 30.0 days?
1
Expert's answer
2011-06-22T13:47:12-0400
The power dissipated by a linear resistor may be calculated from its resistance, and voltage or current. The formula is a combination of Ohm's law and Joule's laws:
P = V2/R = 1102/104 = 12100/10000 = 12.1 W.
Working 30 days we will have
Ptot = 12.1*30*24 = 8712.
Finally we get
$ = 8712*0.05 = 435.6.